OpenStudy (anonymous):

Use completing the square to graph 4x^2+8x+3y^2-18y=-15

OpenStudy (anonymous):

Divide by 3 for both sides.

OpenStudy (anonymous):

and then complete the square for y only first.

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 }x^2+\frac{ 8 }{ 3 }x+y^2-6y=-5\]

OpenStudy (anonymous):

\[y^2-6y=-5\] Ignore the x's for a moment and complete the square for this first.

OpenStudy (anonymous):

y=5 and y=1

OpenStudy (anonymous):

No...complete the square without finding the y values.

OpenStudy (anonymous):

just find it in this form: \[(y-a)^2=b+a^2\]

OpenStudy (anonymous):

\[(y-3)^{2}=4\]

OpenStudy (anonymous):

Put it in this form sorry. not find.

OpenStudy (anonymous):

okay you just did the first part. Now. \[\frac{ 4 }{ 3 }x^2+\frac{ 8 }{ 3 }x+(y-3)^2=4\]

OpenStudy (anonymous):

Now multiply by 3 to both sides.

OpenStudy (anonymous):

\[4x^2+8x+3(y-3)^2=12\]

OpenStudy (anonymous):

Now to complete the square the equation of the x's must be monic meaning it has to have no coefficient.

OpenStudy (anonymous):

That means we have to divide by 4 now to get rid of the coefficient of x^2

OpenStudy (anonymous):

You will get this result when you divide both sides by 4. \[x^2+2x+\frac{3}{4}(y-3)^{2}=3\]

OpenStudy (anonymous):

No don't worry about the \["+\frac{3}{4}(y-3)^2"\]

OpenStudy (anonymous):

Complete the square for x now as usual. \[x^2+2x=3\]

OpenStudy (anonymous):

@imnotatowel

OpenStudy (anonymous):

\[(x+1)^{2}=4\]

OpenStudy (anonymous):

Good. Now you combine the result and you get: \[(x+1)^2+\frac{3}{4}(y-3)^2=4\]

OpenStudy (anonymous):

multiply by 4 to both sides just to get rid of that fraction in the equation.

OpenStudy (anonymous):

wait don't multiply just yet.

OpenStudy (anonymous):

move the \[\frac{3}{4}(y-3)^2\] to the RHS.

OpenStudy (anonymous):

RHS?

OpenStudy (anonymous):

right hand side?

OpenStudy (anonymous):

Wait. Nah, just multiply by 4 to both sides.

OpenStudy (anonymous):

Forget about moving it the Right hand side,

OpenStudy (anonymous):

moving it to*

OpenStudy (anonymous):

Now you know how to graph that? If you don't know how to graph that, then just plot points and see where that leads you.

OpenStudy (anonymous):

I know x+1 will shift it to the left one, but i don't know what the four does to that

OpenStudy (anonymous):

It's a circle.

OpenStudy (anonymous):

\[x^2+y^2=r^2\] where r is the radius.

OpenStudy (anonymous):

with a radius of 4

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] where (h, k) is the vertex.And r^2=4. r is the radius. so r is the square root of 4.

OpenStudy (anonymous):

centre not vertex*

OpenStudy (anonymous):

the centre*

OpenStudy (anonymous):

How do i plot the centre?

OpenStudy (anonymous):

Wait I figured it out, thanks

OpenStudy (anonymous):

I'm not sure about the 4 and 3 beside the brackets. I might need to ask someone about that part. @AravindG Do you know how to draw this circle? \[4(x+1)^2+3(y-3)^2=12\] And sorry, the radius is sqrt(12) not 2. You forgot to multiply by 4 to boths sides.

OpenStudy (anonymous):

both sides*

OpenStudy (anonymous):

You sure you figured it out @imnotatowel ?

OpenStudy (aravindg):

for the equation you wrote above we can divide the whole equation by 12 to get the form \[(x-a)^2+(y-b)^2=r^2\]

OpenStudy (anonymous):

no wouldn't it be fraction+fraction=1?

OpenStudy (anonymous):

because if you divide by 12 to both sides, you get 1/3(x+1)^2 + 1/4(y-3)^2=1

OpenStudy (aravindg):

oh yeah....gimme a minute

OpenStudy (aravindg):

well actually the figure will be an ellipse!!

OpenStudy (anonymous):

Ah okay, I had an inkling about the numbers beside the brackets. I'm not up to ellipses yet but then again, this isn't my question so, thanks for helping @imnotatowel

OpenStudy (aravindg):

infact general form of an ellipse is \[\large \dfrac{(x-h)^2}{a^2}+\dfrac{(y-b)^2}{b^2}=1\] where (h,k) is centre of the ellipse.Next time ensure you dont get that inky feeling again @Azteck :)

OpenStudy (aravindg):

*inkling

OpenStudy (anonymous):

lolz. Ahaha you can count on it.